# Boosting Scientific Error-Bounded Lossy Compression through Optimized Synergistic Lossy-Lossless Orchestration

### v1

@@ 5.2.3. The design details of cuSZ-Hi lossless pipelines
Figure 7 illustrates the two selective lossless pipelines employed in cuSZ-Hi (compressing quantization codes with Huffman-RRE4-TCMS8-RZE1 pipeline or TCMS1-BIT1-RRE1 pipeline). The numbers in the module names indicate the per-symbol byte widths. In the Huffman-RRE4-TCMS8-RZE1 pipeline, after the Huffman encoding, RRE4 introduces a bitmap to mark (set to 0) and eliminate symbols identical to their predecessors, subsequently compressing the bitmap recursively. Following this, TCMS8 employs a reversible bitwise operation (word << 1) ^ (word >> 63) to convert symbols from two’s complement to magnitude-sign representation. Finally, RZE1 compresses the symbols in a manner similar to RRE4 but marks symbols equal to zero. The substantial clustering zeros produced by Huffman encoding, combined with the integration of two reducing stages, maximize the compression ratio of this pipeline.
@@ 6. Evaluations
In this section, we propose our evaluations of cuSZ-Hi on 6 real-world scientific datasets, together with multiple baselines of state-of-the-art GPU-based scientific lossy compressors. From the results, analysis, and discussion, we can clearly find that cuSZ-Hi achieves optimized compression ratio and quality among all existing works.
@@ 6.1.1. Evaluation platforms
The evaluations are conducted on 2 platforms: one is NERSC-Perlmutter (tes, 2021), facilitated with NVIDIA A100 GPUs. The other is a lab-own workstation with NVIDIA 6000 Ada GPUs. The detailed information on them is in Table 2.
@@ 6.1.2. Evaluation baselines and configurations
We include the following GPU-based scientific lossy compressors in the evaluation as baselines: cuSZ (Tian et al., 2020, 2021; Liu et al., 2024c), cuSZp2 (Huang et al., 2024a), cuZFP (Lindstrom, 2014; cuZFP, 2020), and FZGPU (Zhang et al., 2023a). All of them have been widely discussed by the scientific data compression community. Other existing baselines, such as cuSZx (Yu et al., 2022) and MGARD-GPU (Gong et al., 2023) are excluded from our evaluation due to significant performance drawbacks shown in past evaluations (Huang et al., 2024a; Zhang et al., 2023a; Liu et al., 2024c). To all-roundly present the compression of cuSZ, its three modes are evaluated, including cuSZ-L (Lorenzo predictor + Huffman encoder), cuSZ-I (interpolation predictor + Huffman encoder), and cuSZ-IB (interpolation predictor + Huffman encoder + NVIDIA Bitcomp). Due to the proprietary and architecture-specific nature of NVIDIA Bitcomp, we evaluate its usage as a separate baseline to compare cuSZ-Hi with both full-open-source and proprietary solutions. cuSZp2 has 2 modes. To maximize its compression ratio in evaluation, we mostly apply its outlier mode but will fall back to the more stable plain mode on use cases when the outlier mode compression exhibits errors. For our cuSZ-Hi, as previously introduced in Section 5.2, we evaluate two modes of it with different lossless pipelines: cuSZ-Hi-CR and cuSZ-Hi-TP.
@@ 6.1.3. Evaluation datasets
cuSZ-Hi and the baselines are evaluated on 6 real-world scientific datasets (detailed in Table 3). Those datasets are from diverse scientific domains and sources (Zhao et al., 2020a; Li et al., 2008; Liang et al., 2023), and are widely adopted as representative scientific error-bounded lossy compression benchmarks (Zhao et al., 2021; Li et al., 2023; Lindstrom, 2014; Liu et al., 2024a, c; Zhao et al., 2020a).
Table 3. Information of the datasets in experiments
@@ 6.1.4. Evaluation metrics
  • Rate-distortion: We plot the compression bit rate and the decompression data PSNR for compressors. The bit rate $b$ is the average of bits in the compressed data for each input element (i.e., 32×32$\times$32 × the reciprocal of CR). PSNR (Tao et al., 2019) measures the difference between the input data and decompressed data via value range and mean-square error, and a higher PSNR is better.
@@
  • Fixed-CR visualization: The visual qualities of the reconstructed data from all the compressors at the same CR.
@@ 6.2.1. Compression ratio assessment
Table 4 presents the compression ratios that cuSZ-Hi (2 modes, the compression ratio (CR) preferred mode and the throughput (TP) preferred mode) and 5 baselines achieved on all the evaluated datasets with different error bounds (cuZFP is not covered because it doesn’t support fixed-error-bound mode). In almost all test cases, cuSZ-Hi exhibits the best compression ratio, mainly with the CR-preferred mode. In certain cases under large error bounds, the CR can be extremely high (over 300), and the Huffman tree itself can be non-negligible overhead in the compressed data so that the TP-preferred mode without Huffman encoding may have a higher compression ratio than the CR-preferred mode. With synthetical design updates on both the interpolation module and lossless pipeline, cuSZ-Hi successfully gains the compression ratio improvements over baselines ($>200\%$ for 6/18 cases, $>50\%$ for 11/18 cases), empowering optimized data size reduction over different error-bound constraints. Additionally, it is worth noticing that, as the state-of-the-art compressor in terms of compression ratio, cuSZ-IB relies on NVIDIA-proprietary components. When only compared with non-proprietary solutions, the open-source cuSZ-Hi can achieve much higher CR improvements, ranging from $113\%$ to $2786\%$ in all test cases.
@@ 6.2.2. Rate-distortion assessment
Besides the compression ratio discussed in Section 6.2.1, in most scientific lossy compression use cases, the decompression data quality is also critical for justifying the compression solution’s usability. Regarding a joint assessment of compression ratio and data quality, we profile the PSNR metric of decompressed data from different compressors and plot their rate-distortion curves in Figure 8. For each dataset, the upper rectangle plot shows a full bitrate range for all baselines, and the lower square plot demonstrates a narrow bitrate range to visualize better the rate-distortion of high-ratio compressor cuSZ-Hi and cuSZ-I(B). Among all datasets, The CR-preferred mode of cuSZ-Hi (cuSZ-Hi-CR) has delivered stable and excellent rate-distortion, achieving the best compression ratio on most data PSNR value. For example, on the Miranda dataset, cuSZ-Hi-CR achieves $\approx$ $140\%/120\%$ compression ratio improvement over the best baseline cuSZ-IB when the decompression data PSNR is around 59/63. For compressing the QMCPack dataset, with the decompression data PSNR fixed around 60/76, cuSZ-Hi-CR achieves $\approx$ $215\%/80\%$ compression ratio improvement over cuSZ-IB. Moreover, The throughput-preferred mode of cuSZ-Hi (cuSZ-Hi-TP) presents excellent rate-distortion as well, exhibiting close decompression data quality to cuSZ-Hi-CR and outperforming the best baseline (cuSZ-IB) in many cases. Given the fact that cuSZ-Hi-TP has a higher throughput than both cuSZ-IB and cuSZ-Hi-CR (detailed in Section 6.2.4), it can become a top compression solution for performance-sensitive use tasks, such as in-time streaming data compression and so on.
Figure 8. Rate distortion assessment: global (top) and local (bottom) for each dataset. The highly compressible (low-bitrate) regions are indicated by white canvases on top and are zoomed out on bottom.
@@ 6.2.3. Visualization assessment
In this subsection, we present several detailed showcases of high-quality cuSZ-Hi compression, visualizing the decompression data from different compressors under the fixed compression ratio. Figure 9 presents 2 series of visualization results on 2 data snapshots (JHTDB-pressure-#2500 and RTM-#3600), including the slice visualization of the original data and multiple decompression results from cuSZ-Hi and baselines (some baselines are omitted from the visualizations as they have very limited compression ratio and/or data quality), with the compression ratios aligned to close values (e.g. $\approx 144$ for JHTDB). In each case, in similar compression ratios, when obvious error artifacts appear in the decompression data from baselines (cuSZ-IB, cuSZ-L, and cuZFP), the decompression data of cuSZ-Hi shows the best visualization quality without significant distortions and artifacts from the original data. With the evaluation of both data PSNR and visual quality, we have all-roundly verified the high compression fidelity of cuSZ-Hi.
@@ 6.2.4. Speed assessment
After verifying the excellent compression ratio and data quality of cuSZ-Hi, we also evaluate its efficiency, proving that cuSZ-Hi exhibits competitive compression throughput among high-ratio GPU-based scientific lossy compressors. After profiling the computing throughput (in terms of GPU kernel speed) of several scientific lossy compressors on 2 GPU platforms (NVIDIA A100 and RTX 6000 Ada), we present the results acquired in Figure 10. As cuZFP, cuSZp2, and FZ-GPU are throughput-oriented compressors with limited compress ratios, it is natural that cuSZ-Hi features relatively lower throughput than them. When we focus the comparison on compressors that can deliver decent compression ratio and quality, we can notice that cuSZ-Hi has comparable (for CR-preferred mode) or better (for TP-preferred mode) throughput than cuSZ-L and cuSZ-I(B). On NVIDIA RTX6000 Ada, attributed to its high FP computational power and cache capacity, cuSZ-Hi-CR shows a similar compression speed and an acceptable $\sim 20\%$ decompression time overhead compared to cuSZ-I(B) when achieving quite higher compression ratios. On NVIDIA A100, the compression/decompression overhead of cuSZ-Hi-CR compared to cuSZ-I(B) is also mostly limited within $\sim 25\%$. Additionally, while the compression ratio and data quality of cuSZ-Hi-TP is still better than any baselines, it exhibits a throughput constantly higher than cuSZ-I(B), even outperforming cuSZ-L on NVIDIA Ada 6000, showing $\approx 20\%\sim 40\%$ speed improvement over cuSZ-L (also up to doubled compression speed of cuSZ-I(B)) on both compression and decompression. In conclusion, the satisfactory throughput of cuSZ-Hi successfully guarantees its usability in high-performance scientific computing and data management tasks. Moreover, users can adaptively select either of its two modes, depending on their preference for optimizing compression ratio or computing throughput.
@@ 6.2.5. Ablation study
At the end of our evaluation, we would like to justify the effectiveness of the newly proposed design components in cuSZ-Hi with a systematic ablation study. To this end, besides the complete framework of cuSZ-Hi and the baseline cuSZ-IB, we also evaluated several incremental frameworks of cuSZ-IB, stacking separate cuSZ-Hi features. Those results are reported in Table 5, which include the compression ratios of the aforementioned compression frameworks on 4 datasets. Specifically, revising the data partition and anchor stride can bring $6\%\sim 60\%$ compression improvement over the original cuSZ-IB, and further reordering the quantization codes can raise the compression ratio by an additional $6\%\sim 55\%$. Moreover, bringing auto-tuned selection between multi-dimensional and one-dimensional interpolation in the compression framework increases the compression ratio for $8\%\sim 19\%$, and finally, the cuSZ-Hi-CR with optimized lossless encoding pipeline gets a compression ratio which is $16\%\sim 50\%$ higher than before. According to Table 5, each design detail described in Section 5 has contributed well to the improvement of cuSZ-Hi compression ratio.
@@ 7. Conclusion
Scientific error-bounded lossy compression is an essential technique for exascale scientific data management. Unfortunately, as fast-evolving GPUs play a more and more important role in scientific supercomputing, existing scientific error-bounded lossy compression toolkits on GPU platforms are still immature and sub-optimal. This paper proposes cuSZ-Hi, the first GPU-based high-ratio, high-quality, and open-source scientific error-bounded lossy compressor. Integrating highly advanced and GPU-customized data interpolation schemes and also fine-tuned lossless encoding pipelines, cuSZ-Hi significantly outperforms state-of-the-art GPU-based scientific lossy compressors. When exhibiting close or better throughput than the state-of-the-art high-ratio GPU-based compressor cuSZ-I that facilitates NVIDIA-proprietary encoding module, cuSZ-Hi can significantly improve its compression ratio by up to $200\%+$ in the same compression task. In the future, we will improve cuSZ-Hi in the following aspects: 1) Reduce the interpolation computing overhead on GPU platforms; 2) Explore and design better lossless encoders for quantization codes; 3) Enhance the flexibility of compression, such as develop an auto-selection mechanism for different data compressor archetypes and/or lossless pipelines to dynamically fit different data characteristics and compression requirements.
8. Acknowledgments
This research was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research (ASCR), under contracts DE-AC02-06CH11357. This work was also supported by the National Science Foundation (Grant Nos. 2104023, 2311875, 2344717).

### v2

@@ 5.2.3. The design details of cuSZ-Hi lossless pipelines
Figure 7 illustrates the two selective lossless pipelines employed in cuSZ-Hi (compressing quantization codes with HF-RRE4-TCMS8-RZE1 pipeline or TCMS1-BIT1-RRE1 pipeline). The numbers in the module names indicate the per-symbol byte widths. In the HF-RRE4-TCMS8 -RZE1 pipeline, after the Huffman encoding, RRE4 introduces a bitmap to mark (set to 0) and eliminate symbols identical to their predecessors, subsequently compressing the bitmap recursively. Following this, TCMS8 employs a reversible bitwise operation
(word << 1) ^ (word >> 63)
to convert symbols from two’s complement to magnitude-sign representation. Finally, RZE1 compresses the symbols in a manner similar to RRE4 but marks symbols equal to zero. The substantial clustering of zeros produced by Huffman encoding, combined with the integration of two reducing stages, maximizes the compression ratio of this pipeline.
@@ 6. Evaluations
In this section, we present our evaluations of cuSZ-Hi on 6 real-world scientific datasets, along with multiple baselines from state-of-the-art GPU-based scientific lossy compressors. From the results, analysis, and discussion, we can clearly find that cuSZ-Hi achieves optimized compression ratio and quality among all existing works.
@@ 6.1.1. Evaluation platforms
The evaluations are conducted on 2 platforms: one is NERSC-Perlmutter (tes, 2021), facilitated with NVIDIA A100 GPUs. The other is a lab-owned workstation with NVIDIA 6000 Ada GPUs. The detailed information on them is in Table 2.
@@ 6.1.2. Evaluation baselines and configurations
We include the following GPU-based scientific lossy compressors in the evaluation as baselines: cuSZ (Tian et al., 2020, 2021; Liu et al., 2024c), cuSZp2 (Huang et al., 2024a), cuZFP (Lindstrom, 2014; cuZFP, 2020), and FZGPU (Zhang et al., 2023a). All of them have been widely discussed by the scientific data compression community. Other existing baselines, such as cuSZx (Yu et al., 2022) and MGARD-GPU (Gong et al., 2023) are excluded from our evaluation due to significant performance drawbacks shown in past evaluations (Huang et al., 2024a; Zhang et al., 2023a; Liu et al., 2024c). To all-roundly present the compression of cuSZ, its three modes are evaluated, including cuSZ-L (Lorenzo predictor + Huffman encoder), cuSZ-I (interpolation predictor + Huffman encoder), and cuSZ-IB (interpolation predictor + Huffman encoder + NVIDIA Bitcomp). Due to the proprietary and architecture-specific nature of NVIDIA Bitcomp, we evaluate its usage as a separate baseline to compare cuSZ-Hi with both full-open-source and proprietary solutions. cuSZp2 has 2 modes. To maximize its compression ratio in evaluation, we mostly apply its “outlier mode” but will fall back to the more stable “plain mode” on use cases when the outlier mode compression exhibits errors. For our cuSZ-Hi, as previously introduced in §5.2, we evaluate two modes of it with different lossless pipelines: cuSZ-Hi-CR and cuSZ-Hi-TP.
@@ 6.1.3. Evaluation datasets
cuSZ-Hi and the baselines are evaluated on 6 real-world scientific datasets (detailed in Table 3). Those datasets originate from diverse scientific domains and sources (Zhao et al., 2020a; Li et al., 2008; Liang et al., 2023), and are widely adopted as representative benchmarks for error-bounded scientific lossy compression (Zhao et al., 2021; Li et al., 2023; Lindstrom, 2014; Liu et al., 2024a, c; Zhao et al., 2020a).
Table 3. Information on the datasets in experiments
@@ 6.1.4. Evaluation metrics
  • Rate-distortion: We plot the compression bit rate and the decompression data PSNR for compressors. The bit rate $b$ is the average of bits in the compressed data for each input element (i.e., $32\times$ the reciprocal of CR). PSNR (Tao et al., 2019) measures the difference between the input data and decompressed data via value range and mean-square error, and a higher PSNR is better.
@@
  • Fixed-CR visualization: The visual qualities of reconstructed data from all the compressors at the same CR.
@@ 6.2.1. Compression ratio assessment
Table 4 presents the compression ratios that cuSZ-Hi (2 modes, the compression ratio (CR) preferred mode and the throughput (TP) preferred mode) and 5 baselines achieved on all the evaluated datasets with different error bounds (cuZFP is not covered because it doesn’t support fixed-error-bound mode). In almost all test cases, cuSZ-Hi exhibits the best compression ratio, mainly with the CR-preferred mode. In certain cases under large error bounds, the CR can be extremely high (over 300), and the Huffman tree itself can be a non-negligible overhead in the compressed data, so that the TP-preferred mode without Huffman encoding may have a higher compression ratio than the CR-preferred mode. With synthetic design updates on both the interpolation module and lossless pipeline, cuSZ-Hi successfully gains the compression ratio improvements over baselines ($200\%$ for 6/18 cases, $50\%$ for 11/18 cases), empowering optimized data size reduction over different error-bound constraints. Additionally, it is worth noticing that, as the state-of-the-art compressor in terms of compression ratio, cuSZ-IB relies on NVIDIA-proprietary components. When compared only with non-proprietary solutions, the open-source cuSZ-Hi can achieve significantly higher CR improvements, ranging from $113\%$ to $2786\%$ in all test cases.
@@ 6.2.2. Rate-distortion assessment
Besides the compression ratio discussed in §6.2.1, in most scientific lossy compression use cases, the decompression data quality is also critical for justifying the usability of the compression solution. Regarding a joint assessment of compression ratio and data quality, we profile the PSNR metric of decompressed data from different compressors and plot their rate-distortion curves in Figure 8. For each dataset, the upper rectangle plot displays the full bitrate range for all baselines, and the lower square plot illustrates a narrow bitrate range to visualize better the rate-distortion of the high-ratio compressor cuSZ-Hi and cuSZ-I(B). Among all datasets, the CR-preferred mode of cuSZ-Hi (cuSZ-Hi-CR) has delivered stable and excellent rate-distortion, achieving the best compression ratio on most data PSNR values. For example, on the Miranda dataset, cuSZ-Hi-CR achieves $\approx$ $140\%/120\%$ compression ratio improvement over the best baseline cuSZ-IB when the decompression data PSNR is around 59/63. For compressing the QMCPack dataset, with the decompression data PSNR fixed around 60/76, cuSZ-Hi-CR achieves $\approx$ $215\%/80\%$ compression ratio improvement over cuSZ-IB. Moreover, the throughput-preferred mode of cuSZ-Hi (cuSZ-Hi-TP) presents excellent rate-distortion as well, exhibiting close decompression data quality to cuSZ-Hi-CR and outperforming the best baseline (cuSZ-IB) in many cases. Given the fact that cuSZ-Hi-TP has a higher throughput than both cuSZ-IB and cuSZ-Hi-CR (detailed in §6.2.4), it can become a top compression solution for performance-sensitive use tasks, such as in-time streaming data compression and so on.
Figure 8. Rate distortion assessment: global (top) and local (bottom) for each dataset. The highly compressible (low-bitrate) regions are indicated by white canvases on top and are zoomed out on the bottom.
@@ 6.2.3. Visualization assessment
Here, we present several detailed showcases of high-quality cuSZ-Hi compression, visualizing the decompressed data from different compressors at a fixed compression ratio. Figure 9 presents 2 series of visualization results on 2 data snapshots (JHTDB-pressure-#2500 and RTM-#3600), including the slice visualization of the original data and multiple decompression results from cuSZ-Hi and baselines (some baselines are omitted from the visualizations as they have very limited compression ratio and/or data quality), with the compression ratios aligned to close values (e.g. $\approx 144$ for JHTDB). In each case, in similar compression ratios, when highly noticeable error artifacts appear in the decompression data from baselines (cuSZ-IB, cuSZ-L, and cuZFP), the decompression data of cuSZ-Hi shows the best visualization quality without significant distortions and artifacts from the original data. With the evaluation of both data PSNR and visual quality, we have all-aroundly verified the high compression fidelity of cuSZ-Hi.
@@ 6.2.4. Speed assessment
After verifying the excellent compression ratio and data quality of cuSZ-Hi, we evaluate its efficiency, proving that cuSZ-Hi exhibits competitive compression throughput among GPU-based scientific lossy compressors. After profiling the computing throughput (in terms of GPU kernel speed) of several scientific lossy compressors on 2 GPU platforms (NVIDIA A100 and RTX 6000 Ada), we present the results acquired in Figure 10. As cuZFP, cuSZp2, and FZ-GPU are throughput-oriented compressors with limited compression ratios, it is natural that cuSZ-Hi features relatively lower throughput than them. When we focus the comparison on compressors that can deliver decent compression ratio and quality, we can notice that cuSZ-Hi has comparable (for CR-preferred mode) or better (for TP-preferred mode) throughput than cuSZ-L and cuSZ-I(B). On the NVIDIA RTX 6000 Ada, attributed to its high FP computational power and cache capacity, cuSZ-Hi-CR shows a similar compression speed and an acceptable $\sim 20\%$ decompression time overhead compared to cuSZ-I(B) when achieving relatively higher compression ratios. On NVIDIA A100, the compression/decompression overhead of cuSZ-Hi-CR compared to cuSZ-I(B) is also mostly limited within $\sim 25\%$. Additionally, while the compression ratio and data quality of cuSZ-Hi-TP is still better than any baselines, it exhibits a throughput constantly higher than cuSZ-I(B), even outperforming cuSZ-L on NVIDIA Ada 6000, showing $\approx 20\%\sim 40\%$ speed improvement over cuSZ-L (also up to doubled compression speed of cuSZ-I(B)) on both compression and decompression. In conclusion, the satisfactory throughput of cuSZ-Hi successfully guarantees its usability in high-performance scientific computing and data management tasks. Moreover, users can adaptively select either of its two modes, depending on their preference for optimizing compression ratio or computing throughput.
@@ 6.2.5. Ablation study
At the end of our evaluation, we justify the effectiveness of the newly proposed design components in cuSZ-Hi with a systematic ablation study. Besides the complete framework of cuSZ-Hi and the baseline cuSZ-IB, we also evaluated several incremental frameworks of cuSZ-IB, stacking separate cuSZ-Hi features. Those results are reported in Table 5, which includes the compression ratios of the aforementioned compression frameworks on 4 datasets. Specifically, revising the data partition and anchor stride can achieve a compression improvement of $6\%\sim 60\%$ over the original cuSZ-IB, and further reordering the quantization codes can increase the compression ratio by an additional $6\%\sim 55\%$. Moreover, incorporating auto-tuned selection between multi-dimensional and one-dimensional interpolation into the compression framework increases the compression ratio by $8\%\sim 19\%$. Ultimately, the cuSZ-Hi-CR with an optimized lossless encoding pipeline achieves a compression ratio that is $16\%\sim 50\%$ higher than before. According to Table 5, each design detail described in §5 has contributed well to the improvement of cuSZ-Hi compression ratio.
@@ 7. Conclusion
Scientific error-bounded lossy compression is an essential technique for exascale scientific data management. Unfortunately, as fast-evolving GPUs play an increasingly important role in scientific supercomputing, existing scientific error-bounded lossy compression toolkits on GPU platforms remain immature and suboptimal. This paper proposes cuSZ-Hi, the first GPU-based high-ratio, high-quality, and open-source scientific error-bounded lossy compressor. Integrating highly advanced and GPU-customized data interpolation schemes and also fine-tuned lossless encoding pipelines, cuSZ-Hi significantly outperforms state-of-the-art GPU-based scientific lossy compressors. When exhibiting close to or better throughput than the state-of-the-art high-ratio GPU-based compressor cuSZ-I, which utilizes an NVIDIA-proprietary encoding module, cuSZ-Hi can significantly improve its compression ratio by up to $200\%$ in the same compression task. In the future, we will improve cuSZ-Hi in the following aspects: 1) Reduce the interpolation computing overhead on GPU platforms; 2) Explore and design better lossless encoders for quantization codes; 3) Enhance the flexibility of compression, such as develop an auto-selection mechanism for different data compressor archetypes and/or lossless pipelines to fit different data characteristics and compression requirements dynamically.
Acknowledgements.
This research was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research (ASCR), under contracts DE-AC02-06CH11357. This work was also supported by the National Science Foundation (Grant Nos. 2104023, 2311875, 2344717, 2514034, 2514035, and 2514036).